Data archive is back. There may be a little lag in getting data from the last few nights.
They are installing new power outlets in our server room. The computers, including mnem, will be down until 10/9.
The MAWS is back and working fine. The serial port the MAWS was connected to showed signs of intermittent failure. A backup card was installed, and the port changed (ttyS0 -> ttyS4). The opportunity was taken to re-evacuate the dewar, which was not cooling completely. The telescope is on and functioning.
The vaisala weather station is not responding. The telescope will be down for at least a few days until the station can be inspected.
I’ve updated the RCT webpage to highlight science. I’ve moved the telescope maintenance issues and bugs to a separate location on the site, but you will still get posts will still get to the RSS. Please review and let me know if have suggestions or comments.
SII filters (6717, and 6730) are now in filter positions 15 and 16, replacing the UV_continuum and OH filters. The filters now available are:
This information can also be found at RCT Filters
I thought that it might be nice if there was a place where I could regularly post about the happenings at the telescope and anything else RCT-related.
My aim is to post in this blog at least once a week to keep everyone updated on how the telescope is performing.
The type of items I will be posting on:
- Current telescope stats (good observing nights, images taken etc)
- Any observing issues
- Pretty pictures
Please send me your feedback – I want to make this blog as useful as possible!